How do I substitute a problem like this?
y = x + 3
y = 2x + 12

Found 3 solutions by mananth, ikleyn, timofer:
Answer by mananth(16949)

y = x + 3
y = 2x + 12
SO
x+3 = 2x+12
2x-x=3-12
x=-9
plug value of x in 1
y=x+3
y=-9+3
y=-6
(-6,-9)is the ordered solution

Answer by ikleyn(53570)

.
How do I substitute a problem like this?
y = x + 3
y = 2x + 12
~~~~~~~~~~~~~~~~~~~~~~~~


In his post, @mananth obtained the solution x = -9, y = -6.

This is correct.

But then, @mananth points that the ordered solution is (-6,-9).

This is NOT an ordinary/commonly accepted order.

The ordinary/commonly accepted order is opposite (x,y) = (-9,-6).



Answer by timofer(139)

The two expressions for y are equal. How do you substitute? Simple!